OpenStudy (anonymous):

a rock is dropped from a sea cliff and the sound of it striking the ocean is heard 3.2s later if the speed of sound is 340m/s how high is the cliff

OpenStudy (anonymous):

Because the rock is falling straight down we just need to multiply the speed at witch the sound came back into the time it took to come back. d=s * t and we get an answer of d= 340m/s * 3.2 = 1088m if you want to find the time it took; The distance covered by a falling object can be found by the equation s=1/2gt^2 and we can re arrange this equation to solve for time; t= square root{(1/2>)/s} and we substitute our distance into this equation.